OpenStudy (anonymous):

Choose the correct simplification of (3x^3y^4z^4)(2x^3y^4z^2). Answer 6x^6y^8z^6 5x^9y^16z^8 5x^6y^8z^6 6x^9y^16z^8

OpenStudy (anonymous):

3r option

OpenStudy (anonymous):

3rd*

OpenStudy (accessdenied):

Hmm.. well, look at the first coefficients; we have a 2 multiplied to a 3. 2*3=6, so #3 couldn't work

OpenStudy (anonymous):

oh shiz i see we have to multiply

OpenStudy (anonymous):

last option

OpenStudy (accessdenied):

well, now if we multiply x^3 * x^3, we have the same base so essentially we just add exponents: x^(3+3)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

its the last option though right?

OpenStudy (accessdenied):

x^(3+3) = x^6, but there's no x^6 in the last option... so it can't be that. :P

OpenStudy (anonymous):

first 1?

OpenStudy (accessdenied):

Yes. :) In general, if we have: \(\large x^a \times x^b\) Then it is the same as \(\large x^{a+b}\)

OpenStudy (anonymous):

k thanks :D